Discrete Mathematics:

Assume you have a list of numbers, A, and A[i] represents the ith element of the list. Now please define a recursive function, f(n), which represents the sum of the first n elements in the list for n>=0.

Hint 1: a recursive function needs defining its initial value and the relationship between neighbors.

Hint 2: f(n) = A[1] + A[2] + A[3] + … + A[n-1] + A[n] is not what I am asking here.)
Answer #1

Base case:

f(0) = 0                          //because no elements will be added n = 0

Recursive relation:

f(n) = f(n-1) + A(n)

Ex: suppose A[1] = 5, from above relation we have to get f(1) = 5. Let's see

f(n) = f(n-1) + A[n]

f(1) = f(0) + A[1]

       = 0 + 5                              {f(0) = 0 // base case}

       = 5

similarly we can see for n = 2,3....


Note: If you have any doubts please comment.

It will be great help If you like.